# HyperMono: A Monotonicity-aware Approach to Hyper-Relational Knowledge Representation

### v1

@@ Abstract
In a hyper-relational knowledge graph (HKG), each fact is composed of a main triple associated with attribute-value qualifiers, which express additional factual knowledge. The hyper-relational knowledge graph completion (HKGC) task aims at inferring plausible missing links in a HKG. Most existing approaches to HKGC focus on enhancing the communication between qualifier pairs and main triples, while overlooking two important properties that emerge from the monotonicity of the hyper-relational graphs representation regime. Stage Reasoning allows for a two-step reasoning process, facilitating the integration of coarse-grained inference results derived solely from main triples and fine-grained inference results obtained from hyper-relational facts with qualifiers. In the initial stage, coarse-grained results provide an upper bound for correct predictions, which are subsequently refined in the fine-grained step. More generally, Qualifier Monotonicity implies that by attaching more qualifier pairs to a main triple, we may only narrow down the answer set, but never enlarge it. This paper proposes the HyperMono model for hyper-relational knowledge graph completion, which realizes stage reasoning and qualifier monotonicity. To implement qualifier monotonicity HyperMono resorts to cone embeddings. Experiments on three real-world datasets with three different scenario conditions demonstrate the strong performance of HyperMono when compared to the SoTA.
1 Introduction
@@
Knowledge graphs (KGs) store graph-like knowledge as a collection of factual triples [1, 2, 3]. Each triple $(h,r,t)$ in a KG represents a connection between the head entity h and the tail entity t through the relation type $r$, e.g., (James Harden, member_of_team, Philadelphia 76ers) in Figure 1. Knowledge graphs provide the backbone of various applications like query answering [4, 5, 6], search [7, 8], entity typing [9, 10] and recommendation systems [11, 12]. In hyper-relational knowledge graphs (HKGs), hyper-relational facts enhance binary relational facts by augmenting triples with supplementary knowledge expressed in the form of attribute-value qualifier pairs $(a:v)$. For example, in Figure 1, the triple (James Harden, member_of_team, Philadelphia 76ers) is contextualized by the qualifier pairs {(start_time: 2019), (end_time: 2023), (teammate: P.J. Tucker), (part_of: Atlantic Division)}, describing teams of the Atlantic Division in which James Harden and P.J. Tucker played together from 2019 to 2023. Like standard KGs, HKGs also suffer from incompleteness. The hyper-relational knowledge graph completion (HKGC) task, which aims at inferring missing links in HKGs, has therefore become an important research topic.
@@
Figure 1: Qualifier Monotonicity: The box on the right presents all five answers to the query $q=$ (James Harden, member_of_team, ?). If qualifiers are added to the query, the number of possible answers is reduced, e.g., if we add the qualifier pairs (start_time, 2019) and (end_time, 2023) to $q$, Oklahoma City Thunder is not a possible answer anymore.
@@
A wide variety of approaches to HKGC have been already proposed, including embedding-based [13, 14, 15, 16], transformer-based with graph neural networks (GNNs) [17, 18] and transformer-based without GNNs methods [19, 20, 21, 22, 23]. Despite the substantial progress achieved so far, each of these approaches has its own disadvantages: embedding-based methods can be categorized into n-ary, key-value pairs and main triple+key-value pairs depending on how facts are defined. While hyper-relational knowledge is effectively represented, the modeling of distinct auxiliary information is done independently. This independence imposes constraints on the interaction between key-value pairs within a given fact. Transformer-based with GNNs methods mainly rely on multi-layer GNNs operations to encode qualifier-pairs information into the main triple, inevitably introducing noise that propagates as the number of GNN layers increases. In addition, the use of GNNs introduces a considerable computational burden. Transformer-based without GNNs methods capture the structure of HKGs within the transformer architecture, thereby avoiding the increase in computational complexity. However, although these methods are in a dominant position on the HKGC task, they overlook two important properties: Two-Stage Reasoning and Qualifier Monotonicity. When inferring missing entities, two-stage reasoning permits to first reason about binary relational knowledge and then, guided by the results from the previous step, reason about hyper-relational knowledge. Qualifier monotonicity implies that the addition of qualifier information to triples might allow for additional inferences, without invalidating those derived solely from considering main triples. Specifically, these two properties can be described as follow.
  1. 1.
@@
  2. 2.
    Qualifier Monotonicity generally implies that for a given hyper-relational query $q$, as the number of qualifier pairs in $q$ increases, the answer set of $q$ over a HKG might shrink but it never expands, while conversely removing qualifiers from $q$ could only lead to more possible answers. For example, for the query (James Harden, member_of_team, ?) in Figure 1, ? has five possible answers {Oklahoma City Thunder, Houston Rockets, Brooklyn Nets, Philadelphia 76ers, Los Angeles Clippers}. If the qualifier pairs (start_time: 2019) and (end_time: 2023) are added, then Oklahoma City Thunder is not a possible answer anymore. The main challenge for qualifier monotonicity lies in devising a representation that accurately captures its essence. It should posses the capability to faithfully capture the shrinking-like dynamics induced by the incorporation of qualifiers. In the context of embedding-based methods, ShrinkE [16] has explored qualifier monotonicity through box embeddings. However, the inherent limitations of the embedding-based methods lead to a large gap between ShrinkE’s accuracy and that of SoTA models, as shown in § 5.1, Table II.
@@
  • We conduct experiments on three real-world datasets under three different conditions. The obtained results demonstrate the strong performance when compared with SoTA models.
2 Related Work
In this section, we analyze the existing work in the field of HKGC. They can be categorized as follows: Embedding-based Methods (§ 2.1), Transformer-based with GNNs Methods (§ 2.2), Transformer-based without GNNs Methods (§ 2.3), and Other Methods (§ 2.4).
2.1 Embedding-based Methods
@@
Depending on how the facts are represented, embedding-based methods can be classified as n-ary, key-value pairs and main triple+key-value pairs. (i) In n-ary methods, each fact consists of a pre-defined relation r with n corresponding values, denoted as $r(v_{1},v_{2},...,v_{n})$. For instance, m-TransH [13] and RAE [14] generalize TransH [25] by projecting all entities onto a relation-specific hyperplane, while m-CP [26] generalizes Canonical Polyadic (CP) decomposition [27], n-TuckER [28] and GETD [28] generalize TuckER [29] to the higher-arity case. However, these methods convert n-ary facts using one abstract relation, ignoring the semantic distinction of different values for each relation. (ii) In key-value pairs methods, each fact is viewed as a set of attribute-value pairs, denoted as $\{(a_{i}:v_{i})\}_{i=1}^{m}$. For instance, NaLP [15] employs a convolutional neural network followed by a multi-layer perceptron to measure the validity of a fact, RAM [30] further encourages the relatedness among different attributes and between an attribute and all corresponding values. However, these methods treat all attribute-value pairs equally and do not distinguish the main triple from the complete fact. (iii) Main triple+key-value pairs methods regard each fact as a main triple with attribute-value pairs, denoted as $((h,r,t),\{(a_{i}:v_{i})\}_{i=1}^{m})$. For example, HINGE [15] and NeuInfer [31] independently consider each auxiliary pair in the main triple. MSeaKG [32] develops a generic message passing function to encode hyper-relational facts. However, as different auxiliary information is modeled independently, these methods do not properly encode the interactions between attribute-value pairs within fact, further leads to a certain degree of semantic gap between different attribute pairs.
2.2 Transformer-based with GNNs Methods
@@
StarE [17] adopts a message passing based graph encoder CompGCN [33] to obtain the entity and relation embeddings, which are further passed through a transformer decoder to obtain a probability distribution over all entities. QUAD [18] simultaneously takes into account the communication from qualifiers to the main triple and the flow of information from the main triple to qualifiers. It also introduces an auxiliary task to further improve performance. However, these methods primarily employ multi-layer GNNs to incorporate hyper-relational knowledge into the main triple, which inevitably introduces two problems. On the one hand, although multi-layers GNNs operations can encode entity-neighborhood information, they may introduce noise, which can be propagated in a cascading fashion as the number of layers increases. On the other hand, GNNs operations incur in an additional time and space overhead [34, 35, 36], which brings greater challenges to efficiently train these models.
2.3 Transformer-based without GNNs Methods
@@
Many methods abandon the use of GNNs operations and instead incorporate hyper-relational knowledge through various transformer structures. For example, HyTrans [19] extends StarE by substituting the graph encoder with a light-weight relation or entity embeddings. GRAN [20] implements edge-biased fully-connected attention into transformers to encode the graph structure and its heterogeneity. HyNT [21] defines a specialized context transformer to exchange information between the main triple and its qualifiers, HAHE [22] proposes a hierarchical attention structure with global-level and local-level attention to simultaneously model graph-like and sequential content. HyperFormer [23] introduces a mixture-of-experts strategy into the feed-forward layers of a transformer to strengthen its representation capabilities. However, these methods have two shortcomings: on the one hand, they only consider a fine-grained inference steps with hyper-relational knowledge and overlook coarse-grained inferences solely based on the main triples. On the other hand, they do not explicitly embody qualifier monotonicity. We note in passing that although the embeeding-based method ShrinkE [16] takes qualifier monotonicity into account, its embedding-based nature results in a large gap in performance when compared with the state-of-the-art, as shown in § 5.1, Table II.
2.4 Other Methods
Some researchers have tried to introduce external knowledge (e.g., schema information) into the representation of entities and relations to assist HKGC tasks. sHINGE [37] captures not only the primary structural information in the triples and associated key-value pairs, but also the schema information encoded by entity-typed triples and corresponding entity-typed key-value pairs. DHGE [38] proposes a dual-view hyper-relational structure that contains a hyper-relational instance view for entities and a hyper-relational ontology view for concepts to encode hierarchical knowledge. However, these methods require acquisition of external knowledge relative to the entity, which poses a greater challenge to their general applicability. Considering the fairness of the experimental comparison and that the main goal is the construction of a more general model, we did not consider these baselines.
3 Background
In this section, we introduce necessary background on Hyper-relational Knowledge Graphs (§ 3.1), Qualifier Monotonicity (§ 3.2), and Cone Embeddings (§ 3.3).
3.1 Hyper-relational Knowledge Graph
@@ 3.2 Qualifier Monotonicity
3.3 Cone Embedding
We present here only necessary background on cone embeddings, for a full discussion we refer the interested reader to [4, 6, 24]. A d-dimensional sector-cone can be parameterized by $\mathcal{C}^{d}(\alpha,\beta)=((\alpha^{1},\beta^{1}),...,(\alpha^{d},\beta^{d}))$, where $\alpha^{i}\in[-\pi,\pi)$ represents the angle between the symmetry axis of the i-th-dimension sector-cone and the positive x axis, and $\beta^{i}$ represents the aperture of the i-th-dimension sector-cone. Figure 3(a) shows a typical 1-dimensional sector-cone. We omit the superscript of $\mathcal{C}^{d}$ if it is clear from context. A variety of operations on cone embeddings can be defined, here we introduce the projection $\mathcal{P}$ and intersection $\mathcal{I}$ operations:
  1. 1.
    Projection Operator $\mathcal{P}$. $\mathcal{C}_{2}$=$\mathcal{P}$($\mathcal{C}_{1}$, $\mathcal{C}_{r}$) computes the projection from an input cone $\mathcal{C}_{1}$ as a head entity set to the set of tail entities $\mathcal{C}_{2}$ via the specific relation cone $\mathcal{C}_{r}$, see Figure 3(b). Following [4], we employ the projection operation to convert from one sector-cone to another sector-cone.
  2. 2.
    Intersection Operator $\mathcal{I}$. $\mathcal{I}$($\mathcal{C}_{1}$, $\mathcal{C}_{2}$) computes the intersection of the set of entities $\mathcal{C}_{1}$ and the set of entities $\mathcal{C}_{2}$, see Figure 3(c). Considering that the intersection operation is universal in a sector-cone space (and that it is not our innovation), we directly follow the specifications defined by [4].
4 Method
Figure 2: HyperMono’s model structure, containing two modules: Head Neighborhood Encoder (HNE) and Missing Entity Predictor (MEP). HNE module contains two sub-modules: Coarse-grained Neighborhood Aggregator (CNA) and Fine-grained Neighborhood Aggregator (FNA). MEP module contains two sub-modules: Triple-based Predictor (TP) and Qualifier-Monotonicity-aware Predictor (QMP).
In this section, we present the two main components of HyperMono’s architecture: HNE: Head Neighborhood Encoder (§ 4.1) and MEP: Missing Entity Predictor (§ 4.2). In the HNE module, we consider two types of relational contexts, and construct the sub-modules Coarse-grained Neighbor Aggregator and Fine-grained Neighbor Aggregator. In the MEP module, we introduce the Triple-based Predictor and Qualifier-Monotonicity-aware Predictor sub-modules to faithfully capture the two-stage reasoning and qualifier monotonicity properties.
4.1 HNE: Head Neighborhood Encoder
@@
To adequately encode the content of the neighbors of a head entity, as a first step, we introduce the Head Neighborhood Encoder (HNE) module, cf.  Figure 2(b). The HNE module considers two types of relational contexts, characterized by the use of either main triples or complete hyper-relational facts to ascertain the neighbors of a given entity: Coarse-grained Neighbor Aggregator (CNA), only considering main triples in $\mathcal{G}_{triple}$; Fine-grained Neighbor Aggregator (FNA), considering hyper-relational facts in $\mathcal{G}$. The resulting neighbor-aware embeddings from the CNA and FNA modules will be used in the subsequent module (§4.2) to provide further evidence, based on the neighbors of $h$, for the prediction of the ???? entity.
@@
  1. 1.
    Neighbor-aware Embeddings. Let $N_{h}=\{(h,r^{\prime},t^{\prime})~{}|~{}$ $(h,r^{\prime},t^{\prime})\in\mathcal{G}_{\textit{triple}}\}$ be the set of neighbors of $h$. For a triple $\textit{tr}=(h,r^{\prime},t^{\prime})\in N_{h}$, we use the placeholder [mask] in place of $h$ to obtain the neighbor sequence $\{[\texttt{mask}],r^{\prime},t^{\prime}\}$. Then, we apply a standard transformer encoder [39] to obtain the [mask] token embedding $\textbf{{N}}_{\textit{tr}}^{h}\in\mathbb{R}^{1\times d}$, $d$ represents the embedding dimension. Clearly, if $n=|N_{h}|$, n different [mask] representations will be obtained. So, we sum up and average all [mask] representations to obtain an aggregated embedding $\textbf{{N}}^{h}$ considering all neighbors of h: $\textbf{{N}}^{h}=mean(\sum_{tr\in N_{h}}\textbf{{N}}_{\textit{tr}}^{h})$.
  2. 2.
@@
  3. 3.
    Aggregating Inference Results. The LEI and GEI mechanisms will generate multiple entity inference results, to unify them, inspired by  [9, 40], we adopt an exponentially weighted pooling method to generate the final inference result:
    $\textbf{{P}}^{h}=pool(\{\textbf{{P}}_{\textit{tr}_{1}}^{h},\textbf{{P}}_{\textit{tr}_{2}}^{h},...,\textbf{{P}}_{\textit{tr}_{n}}^{h},\textbf{{P}}_{G}^{h}\})=\sum_{i=1}^{n,G}\omega_{i}\textbf{{P}}_{\textit{tr}_{i}}^{h}$ (1)
@@
4.2 MEP: Missing Entity Predictor

### v2

@@ Abstract.
In a hyper-relational knowledge graph (HKG), each fact is composed of a main triple associated with attribute-value qualifiers, which express additional factual knowledge. The hyper-relational knowledge graph completion (HKGC) task aims at inferring plausible missing links in a HKG. Most existing approaches to HKGC focus on enhancing the communication between qualifier pairs and main triples, while overlooking two important properties that emerge from the monotonicity of the hyper-relational graphs representation regime. Stage Reasoning allows for a two-step reasoning process, facilitating the integration of coarse-grained inference results derived solely from main triples and fine-grained inference results obtained from hyper-relational facts with qualifiers. In the initial stage, coarse-grained results provide an upper bound for correct predictions, which are subsequently refined in the fine-grained step. More generally, Qualifier Monotonicity implies that by attaching more qualifier pairs to a main triple, we may only narrow down the answer set, but never enlarge it. This paper proposes the HyperMono model for hyper-relational knowledge graph completion, which considers an initial triple-level reasoning phase to be subsequently refined by a triple+qualifiers reasoning step to infer missing entities. To realize qualifier monotonicity, HyperMono represents main triples as cones and each qualifier is modeled as a cone that shrinks the cone of the main triple to which the qualifier has been added. Moreover, HyperMono leverages the neighborhood context of an entity to semantically strengthen its representation. Experiments on three real-world datasets with three different scenario conditions demonstrate the strong performance of HyperMono when compared to the SoTA.
Knowledge Graph, Hyper-relational Graph, Knowledge Graph Completion
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Knowledge representation and reasoningccs: Information systems Data mining
1. Introduction
@@
Knowledge graphs (KGs) store graph-like knowledge as a collection of factual triples (Bordes et al., 2013; Sun et al., 2019; Li et al., 2022). Each triple $(h,r,t)$ in a KG represents a connection between the head entity h and the tail entity t through the relation type $r$, e.g., (James Harden, member_of_team, Philadelphia 76ers) in Figure 1. Knowledge graphs provide the backbone of various applications like query answering (Zhang et al., 2021; Hu et al., 2022; Nguyen et al., 2023b), search (Nguyen et al., 2019; Gu et al., 2019), and recommendation systems (Yang et al., 2022; Zou et al., 2022). In hyper-relational knowledge graphs (HKGs), hyper-relational facts enhance binary relational facts by augmenting triples with supplementary knowledge expressed in the form of attribute-value qualifier pairs $(a:v)$. For example, in Figure 1, the triple (James Harden, member_of_team, Philadelphia 76ers) is contextualized by the qualifier pairs {(start_time: 2019), (end_time: 2023), (teammate: P.J. Tucker), (part_of: Atlantic Division)}, describing teams of the Atlantic Division in which James Harden and P.J. Tucker played together from 2019 to 2023. Like standard KGs, HKGs also suffer from incompleteness. The hyper-relational knowledge graph completion (HKGC) task, which aims at inferring missing links in HKGs, has therefore become an important research topic.
@@
Figure 1. Qualifier Monotonicity: The box on the right presents all five answers to the query $q=$ (James Harden, member_of_team, ?). If qualifiers are added to the query, the number of possible answers is reduced, e.g., if we add the qualifier pairs (start_time, 2019) and (end_time, 2023) to $q$, Oklahoma City Thunder is not a possible answer anymore.
@@
A wide variety of approaches to HKGC have been already proposed, including embedding-based (Wen et al., 2016; Zhang et al., 2018; Rosso et al., 2020; Xiong et al., 2023), transformer-based with graph neural networks (GNNs) (Galkin et al., 2020; Shomer et al., 2022) and transformer-based without GNNs methods (Yu and Yang, 2021; Wang et al., 2021; Chung et al., 2023; Luo et al., 2023; Hu et al., 2023). Despite the substantial progress achieved so far, each of these approaches has its own disadvantages: embedding-based methods can be categorized into n-ary, key-value pairs and main triple+key-value pairs depending on how facts are defined. While hyper-relational knowledge is effectively represented, the modeling of distinct auxiliary information is done independently. This independence imposes constraints on the mutual interaction between key-value pairs within a given fact. Transformer-based with GNNs methods mainly rely on multi-layer GNNs operations to encode qualifier-pairs information into the main triple, inevitably introducing noise that propagates as the number of GNN layers increases. In addition, the use of GNNs introduces a considerable computational burden. Transformer-based without GNNs methods capture the structure of HKGs within the transformer architecture, thereby avoiding the increase in computational complexity. However, although these methods are in a dominant position on the HKGC task, they overlook two important properties: Two-Stage Reasoning and Qualifier Monotonicity. When inferring missing entities, two-stage reasoning permits to first reason about binary relational knowledge and then, guided by the results from the previous step, reason about hyper-relational knowledge, while qualifier monotonicity implies that adding qualifiers information to triples might allow for additional inferences, without invalidating those derived solely from considering main triples. Specifically, these properties can be described as follows.
  1. (1)
@@
  2. (2)
    Qualifier Monotonicity generally implies that for a given hyper-relational query $q$, as the number of qualifier pairs in $q$ increases, the answer set of $q$ over a HKG might shrink but it never expands, while conversely removing qualifiers from $q$ could only lead to more possible answers. For example, for the query (James Harden, member_of_team, ?) in Figure 1, ? has five possible answers {Oklahoma City Thunder, Houston Rockets, Brooklyn Nets, Philadelphia 76ers, Los Angeles Clippers}. If the qualifier pairs (start_time: 2019) and (end_time: 2023) are added, then Oklahoma City Thunder is not a possible answer anymore. The main challenge for qualifier monotonicity lies in devising a representation that accurately captures its essence. It should posses the capability to faithfully capture the shrinking-like dynamics induced by the incorporation of qualifiers. In the context of embedding-based methods, ShrinkE (Xiong et al., 2023) has explored qualifier monotonicity through box embeddings. However, the inherent limitations of the embedding-based methods lead to a large gap between ShrinkE’s accuracy and that of SoTA models, as shown in § 4.1, Table 3.
@@
  • We conduct experiments on three real-world datasets under three different conditions, and various ablation studies on parameter sensitivity, complexity analysis, and model transferability. The obtained results demonstrate the strong performance when compared with SoTA models.
2. Related Work
@@
Embedding-based. Depending on how the facts are represented, embedding-based methods can be classified as n-ary, key-value pairs and main triple+key-value pairs. (i) In n-ary methods, each fact consists of a pre-defined relation r with n corresponding values, denoted as $r(v_{1},v_{2},...,v_{n})$. For instance, m-TransH (Wen et al., 2016) and RAE (Zhang et al., 2018) generalize TransH (Wang et al., 2014) by projecting all entities onto a relation-specific hyperplane, while m-CP (Fatemi et al., 2020) generalizes Canonical Polyadic (CP) decomposition (Trouillon et al., 2017), n-TuckER (Liu et al., 2020) and GETD (Liu et al., 2020) generalize TuckER (Balazevic et al., 2019) to the higher-arity case. However, these methods convert n-ary facts using one abstract relation, ignoring the semantic distinction of different values for each relation. (ii) In key-value pairs methods, each fact is viewed as a set of attribute-value pairs, denoted as $\{(a_{i}:v_{i})\}_{i=1}^{m}$. For instance, NaLP (Rosso et al., 2020) employs a convolutional neural network followed by a multi-layer perceptron to measure the validity of a fact, RAM (Liu et al., 2021) further encourages the relatedness among different attributes and between an attribute and all corresponding values. However, these methods treat all attribute-value pairs equally and do not distinguish the main triple from the complete fact. (iii) Main triple+key-value pairs methods regard each fact as a main triple with attribute-value pairs, denoted as $((h,r,t),\{(a_{i}:v_{i})\}_{i=1}^{m})$. For example, HINGE (Rosso et al., 2020) and NeuInfer (Guan et al., 2020) independently consider each auxiliary pair in the main triple. However, as different auxiliary information is modeled independently, these methods do not properly encode the interactions between attribute-value pairs within fact.
@@
Transformer-based with GNNs. StarE (Galkin et al., 2020) adopts a message passing based graph encoder CompGCN (Vashishth et al., 2020) to obtain the entity and relation embeddings, which are further passed through a transformer decoder to obtain a probability distribution over all entities. QUAD (Shomer et al., 2022) simultaneously takes into account the communication from qualifiers to the main triple and the flow of information from the main triple to qualifiers. It also introduces an auxiliary task to further improve performance. However, these methods primarily employ multi-layer GNNs to incorporate hyper-relational knowledge into the main triple, which inevitably introduces two problems. On the one hand, although multi-layers GNNs operations can encode entity-neighborhood information, they may introduce noise, which can be propagated in a cascading fashion as the number of layers increases. On the other hand, GNNs operations incur in an additional time and space overhead (Chen et al., 2018; Wu et al., 2019; You et al., 2020).
@@
Transformer-based without GNNs. Many methods abandon the use of GNNs operations and instead incorporate hyper-relational knowledge through various transformer structures. For example, GRAN (Wang et al., 2021) implements edge-biased fully-connected attention into transformers to encode the graph structure and its heterogeneity. HyNT (Chung et al., 2023) defines a specialized context transformer to exchange information between the main triple and its qualifiers, HAHE (Luo et al., 2023) proposes a hierarchical attention structure with global-level and local-level attention to simultaneously model graph-like and sequential content. HyperFormer (Hu et al., 2023) introduces a mixture-of-experts strategy into the feed-forward layers of a transformer to strengthen its representation capabilities. However, these methods have two shortcomings: on the one hand, they only consider a fine-grained inference steps with hyper-relational knowledge and overlook coarse-grained inferences solely based on the main triples. On the other hand, they do not explicitly embody qualifier monotonicity. We note in passing that although the embeeding-based method ShrinkE (Xiong et al., 2023) takes qualifier monotonicity into account, its embedding-based nature results in a large gap in performance when compared with the state-of-the-art, as shown in § 4.1, Table 3.
3. Method
3.1. Preliminaries
@@
Cone Embedding.
We present here only necessary background on cone embeddings, for a full discussion we refer the interested reader to (Zhang et al., 2021; Nguyen et al., 2023c, a). A d-dimensional sector-cone can be parameterized by $\mathcal{C}^{d}(\alpha,\beta)=((\alpha^{1},\beta^{1}),...,(\alpha^{d},\beta^{d}))$, where $\alpha^{i}\in[-\pi,\pi)$ represents the angle between the symmetry axis of the i-th-dimension sector-cone and the positive x axis, and $\beta^{i}$ represents the aperture of the i-th-dimension sector-cone. We omit the superscript of $\mathcal{C}^{d}$ if it is clear from context. A variety of operations on cone embeddings can be defined, here we introduce the projection $\mathcal{P}$ and intersection $\mathcal{I}$ operations:
  1. (1)
    Projection Operator $\mathcal{P}$. $\mathcal{C}_{2}$=$\mathcal{P}$($\mathcal{C}_{1}$, $\mathcal{C}_{r}$) computes the projection from an input cone $\mathcal{C}_{1}$ as a head entity set to the set of tail entities $\mathcal{C}_{2}$ via the specific relation cone $\mathcal{C}_{r}$, see Figure 3(b).
  2. (2)
    Intersection Operator $\mathcal{I}$. $\mathcal{I}$($\mathcal{C}_{1}$, $\mathcal{C}_{2}$) computes the intersection of the set of entities $\mathcal{C}_{1}$ and the set of entities $\mathcal{C}_{2}$, see Figure 3(c).
3.2. Model Architecture
We present in this section the two main components of HyperMono’s architecture: Head Neighborhood Encoder (§ 3.2.1) and Missing Entity Predictor (§ 3.2.2).
3.2.1. HNE: Head Neighborhood Encoder
@@
To adequately encode the content of the neighbors of a head entity, as a first step, we introduce the Head Neighborhood Encoder (HNE) module, cf.  Figure 2(b). The HNE module considers two types of relational contexts, characterized by the use of either main triples or complete hyper-relational facts to ascertain the neighbors of a given entity: Coarse-grained Neighbor Aggregator (CNA), only considering main triples in $\mathcal{G}_{triple}$; Fine-grained Neighbor Aggregator (FNA), considering hyper-relational facts in $\mathcal{G}$. The resulting neighbor-aware embeddings from the CNA and FNA modules will be used in the subsequent module (§3.2.2) to provide further evidence, based on the neighbors of $h$, for the prediction of the ???? entity.
Figure 2. HyperMono’s model structure, containing two modules: Head Neighborhood Encoder (HNE) and Missing Entity Predictor (MEP). HNE module contains two sub-modules: Coarse-grained Neighborhood Aggregator (CNA) and Fine-grained Neighborhood Aggregator (FNA). MEP module contains two sub-modules: Triple-based Predictor (TP) and Qualifier-Monotonicity-aware Predictor (QMP).
@@
  1. (1)
    Neighbor-aware Embeddings. Let $N_{h}=\{(h,r^{\prime},t^{\prime})~{}|~{}$ $(h,r^{\prime},t^{\prime})\in\mathcal{G}_{\textit{triple}}\}$ be the set of neighbors of $h$. For a triple $\textit{tr}=(h,r^{\prime},t^{\prime})\in N_{h}$, we use the placeholder [mask] in place of $h$ to obtain the neighbor sequence $\{[\texttt{mask}],r^{\prime},t^{\prime}\}$. Then, we apply a standard transformer encoder (Vaswani et al., 2017) to obtain the [mask] token embedding $\textbf{{N}}_{\textit{tr}}^{h}\in\mathbb{R}^{1\times d}$, $d$ represents the embedding dimension. Clearly, if $n=|N_{h}|$, n different [mask] representations will be obtained. So, we sum up and average all [mask] representations to obtain an aggregated embedding $\textbf{{N}}^{h}$ considering all neighbors of h: $\textbf{{N}}^{h}=mean(\sum_{tr\in N_{h}}\textbf{{N}}_{\textit{tr}}^{h})$.
  2. (2)
@@
  3. (3)
    Aggregating Inference Results. The LEI and GEI mechanisms will generate multiple entity inference results, to unify them, inspired by  (Hu et al., 2022; Pan et al., 2021), we adopt an exponentially weighted pooling method to generate the final inference result:
    (1) $\textbf{{P}}^{h}=pool(\{\textbf{{P}}_{\textit{tr}_{1}}^{h},\textbf{{P}}_{\textit{tr}_{2}}^{h},...,\textbf{{P}}_{\textit{tr}_{n}}^{h},\textbf{{P}}_{G}^{h}\})=\sum_{i=1}^{n,G}\omega_{i}\textbf{{P}}_{\textit{tr}_{i}}^{h}$
@@
3.2.2. MEP: Missing Entity Predictor